OpenStudy (anonymous):

What is the vertex of the parabola defined by the equation (x − 2)2 = -12(y − 2)? A) (-12, 2) B) (2, 2) C) (6, 2) D) (2, -2)

OpenStudy (vishweshshrimali5):

Vertex of a parabola with equation: \[\large{(y-k)^2 = 4a(x-h)}\] is : (h,k)

OpenStudy (igreen):

Someone asked the same question..the answer is B. https://www.desmos.com/calculator

OpenStudy (anonymous):

Welcome to OpenStudy

OpenStudy (vishweshshrimali5):

So, just put (x-2) = 0 and (y-2) = 0 and you will get your answer.